OpenStudy (anonymous):

Which is an example of the associative property? A. –(18 + 3) = –18 + (–3) B. –17 + 0 = –17 C. (–8 + 10) + 4. = –8 + (10 + 4) D. –5 + 7 = 7 + (–5) Is it c???

OpenStudy (skullpatrol):

$$\Huge (–8 + 10) + 4 = –8 + (10 + 4)$$ $$\Huge (a+b)+c=a+(b+c)$$

OpenStudy (skullpatrol):

yes

OpenStudy (skullpatrol):

c is correct

OpenStudy (anonymous):

Thanks!